Please help me solve this problem: The ratio of the weight of an object on Jupiter to its weight on Earth is 8 to 3. How much would a 100 pound person weigh on Jupiter?
Answer by fractalier(6550)

use a proportion for this...
Jupiter/Earth = 8/3 = x/100
Now cross multiply and get
3x = 800
and
x = 800/3 = 266 2/3 lbs